# The Power of Graph Sparsification in the Continual Release Model

### v1

@@ 4.1.1 Algorithm Description
The pseudocode for our algorithm is given in Algorithm 4.1. Our algorithm works as follows. We set the probability of sampling edges from the stream initially to $1$. For each update $e_{t}$, we sample the update into $S$ with probability $p_{t}$ (Algorithm 4.1). If the sampled update is not $\bot$ (Algorithm 4.1), then we initialize counters for its endpoints if one or both of these counters do not already exist (Algorithm 4.1). A sampled $e_{t}$ that is $\bot$ does not increase the size of $S$ (it is a no-op). Then, we iterate through every other edge in $S$ (Algorithm 4.1) and if $e^{\prime}$ shares an endpoint with $e_{t}$ (Algorithm 4.1), then we increment the counter $c_{e^{\prime}}^{w}$ associated with $e^{\prime}$ (Algorithm 4.1). If the counter exceeds our cutoff $\tilde{\alpha}$ (Algorithm 4.1), then we remove $e^{\prime}$ and its corresponding counters from $S$ (Algorithm 4.1). After performing our edge sampling, we now check using SVT whether $|S|$ (the number of edges in $S$) exceeds $\frac{a_{3}\log^{2}n}{\varepsilon\eta^{2}}$ for some large enough constant $a_{3}>0$ (Algorithm 4.1). If the SVT AboveThreshold query is satisfied, then we halve the probability of sampling (Algorithm 4.1) and resample each edge in $S$ with $1/2$ probability (Algorithm 4.1). Finally, we output a new estimate if $|S|/p$ is greater than our previous estimate by a significantly large amount. Specifically, we check via SVT queries (Algorithm 4.1) whether $|S|$ exceeds $p\cdot(1+\eta)^{j_{t}}$. If so, we increment the counter $j_{t}$ and output our new estimate as $(1+\eta)^{j_{t}}$ (Algorithm 4.1); otherwise, we output our old estimate.
@@ 31                         With probability $\nicefrac{{1}}{{2}}$ remove $e$ from $S$ and delete $c_{e}^{u},c_{e}^{v}$
Algorithm 4.1 Sampling $E_{t}^{*}$
4.1.2 Privacy Guarantee
@@
Lemma 4.2.
Algorithm 4.1 is $\varepsilon$-edge differentially private.
@@
By 4.3, since the sensitivity of the sampled set of edges $S$ is upper bounded by 2. The subsample SVT output is $\varepsilon/2$-DP by Theorem 3.13. Now, observe that the probability $p$ of sampling only changes whenever the subsample SVT (Algorithm 4.1) accepts. Since privacy is preserved after post-processing, this implies the statement in the claim. ∎
@@
By definition, ${\mathcal{P}}$ and ${\mathcal{P}}^{\prime}$ defines the sampling probability for all elements in ${\mathcal{S}}$ and ${\mathcal{S}}^{\prime}$ to be $p_{t}$. Hence, it holds that $\frac{\mathbf{P}[{\mathcal{I}}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})})|{\mathcal{P}}]}{\mathbf{P}[{\mathcal{I}}^{\prime}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})})|{\mathcal{P}}^{\prime}]}=1$. By 4.4, $\frac{\mathbf{P}[{\mathcal{P}}]}{\mathbf{P}[{\mathcal{P}}^{\prime}]}\leq e^{\varepsilon/2}$ and our claim follows. ∎
Let ${\mathcal{Q}}$ be the event that $({\mathcal{I}}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{(T)}),{\mathcal{P}})$ and define ${\mathcal{Q}}^{\prime}$ analogously.
{adjustwidth}
3em0em
Claim 4.6.
Let $\mathbf{a}\in\{\text{``above''},\text{``below''},\text{``abort''}\}^{T}$ be a sequence of the estimate SVT answers in Algorithm 4.1. We have
@@
By 4.3, since the sensitivity of the sampled set of edges $S$ is upper bounded by 2, the estimate SVT output is $\varepsilon/2$-DP by Theorem 3.13. ∎
@@ 4.1.3 Approximation Guarantees
Let $E^{*}_{t}=\max_{t^{\prime}\leq t}(|E_{\tilde{\alpha}}^{t^{\prime}}|)$, we now show that we get a $\left(1+\eta,\frac{\log^{2}(n)}{\varepsilon\eta}\right)$-approximation of $E_{t}^{*}$ with high probability. First, we note that $E^{*}_{t}$ satisfies $M(G^{t})\leq E^{*}_{t}\leq(2+\tilde{\alpha})M(G_{t})$ since $E^{*}_{t}\geq|E_{\tilde{\alpha}}^{t}|$, $M(G_{t^{\prime}})\leq M(G_{t})$, and Lemma 4.7.
@@
Lemma 4.9.
Algorithm 4.1 returns a $\left(1+\eta,\frac{\log^{2}(n)}{\varepsilon\eta}\right)$-approximation of $E_{t}^{*}$ for every $t\in[T]$ with probability at least $1-1/\operatorname*{poly}(n)$.
@@
First, let $\tau=\frac{d\log^{2}(n)}{\varepsilon\eta^{2}}$ for some large enough constant $d>0$ and define level $i_{t}$ (starting with $i_{t}=2$) to be the level that contains $E_{t}^{*}$ if $2^{i_{t}-1}\cdot\tau<E^{*}_{t}\leq 2^{i_{t}}\cdot\tau$. Level $i_{t}=1$ is defined as $0\leq E^{*}_{t}\leq 2\cdot\tau$. Let $p_{t}$ denote the marginal probability that an edge is sampled from $E_{\alpha}^{t}$. Note that this probability is prior to sub-sampling on Algorithm 4.1. In a perfect situation, at time $t$, edge $e$ is sampled in level $i$ with probability $p_{t}^{\prime}:=2^{-i_{t}}$ when $i_{t}\geq 2$ and with probability $p_{t}^{\prime}=1$ when $i_{t}=1$, but this is not the case as our algorithm determines $p_{t}$ adaptively depending on the size of the sampled edges at time $t$.
SVT Budget.
We show that with probability at least $1-1/\operatorname*{poly}(n)$, neither of the SVT “above” budgets $Q_{1}$, $Q_{2}$ in Algorithm 4.1 are exceeded. First, for the sub-sampling SVT, we take $T=\operatorname*{poly}(n)$ noisy comparisons with ${\textsf{Lap}}(O(\varepsilon/\log(n)))$ noise. Hence with probability $1-1/\operatorname*{poly}(n)$, all Laplace realizations are of order $O(\varepsilon^{-1}\log^{2}(n))$.
$Q_{1}=a_{1}\log(n)$ bounds the number of times the subsampling SVT in Algorithm 4.1 is satisfied. Each time this SVT is satisfied, we must have
@@
for sufficiently large constant $a_{3}>0$, leading to the probability of sampling edges being halved. The expected number of edges that are sampled at any point is $p_{t}\cdot\lvert E_{\tilde{\alpha}}^{t}\rvert\leq p_{t}\cdot n^{2}$. Suppose $a_{1}$ is sufficiently large and we halved the sampling probability $\nicefrac{{Q_{1}}}{{2}}\geq\lceil\log(n^{2})\rceil$ times so that $p_{t}\leq\frac{1}{n^{2}}$. We can upper bound the probability that the SVT answers “above” by considering the experiment where we obtain a sample $\tilde{S}_{t}$ by sampling each edge with probability $\tilde{p}=\frac{1}{\lvert E_{\tilde{\alpha}}^{t}\rvert}\geq p_{t}$. By a multiplicative Chernoff bound (Theorem 3.1), for $\psi=\frac{a_{3}}{2}\varepsilon^{-1}\log^{2}(n)-1$,
@@
Concentration.
In a perfect situation, edge $e$ is sampled at time $t$ with probability $p^{\prime}_{t}=2^{-i_{t}}$ for $i_{t}\geq 2$ and with probability $p^{\prime}_{t}=1$ for $i_{t}=1$. In that case, we can use the multiplicative Chernoff bound to bound the probability that our estimate concentrates as our sampling probability would be sufficiently high enough to do so. However, it is not the case that $p_{t}$ is guaranteed to be $2^{-i_{t}}$ since it is determined adaptively by Algorithm 4.1. Hence, we need a slightly more sophisticated analysis than simply using the multiplicative Chernoff bound with $p^{\prime}_{t}$.
We consider two cases. If $p_{t}\geq p^{\prime}_{t}$, then we can lower bound the probability of success by what we would obtain using $p_{t}^{\prime}$ and we can use the multiplicative Chernoff bound in this case. Now, suppose that $p_{t}<p^{\prime}_{t}$; then, we claim that this case never occurs in Algorithm 4.1 with probability at least $1-1/\operatorname*{poly}(n)$. We now formally present these arguments.
If $p_{t}\geq p_{t}^{\prime}$, then we use a multiplicative Chernoff bound (Theorem 3.1) to bound our probability of success. Let $s_{t}$ be the number of edges we sampled for $E_{\tilde{\alpha}}^{t}$ prior to any sub-sampling. That is, we obtain $s_{t}$ by sampling with probability $p_{t}$ before Algorithm 4.1. For $\psi:=\eta\cdot E_{t}^{*}/\lvert E_{\tilde{\alpha}}^{t}\rvert$,
@@
Now, we consider the case when $p_{t}<p_{t}^{\prime}$. We show that with probability at least $1-1/\operatorname*{poly}(n)$, this case does not occur. We prove this via induction on $t$. For $t=1$, $p_{t}=p_{t}^{\prime}$ trivially since both equal $1$. Now, we assume our claim holds for $t$ and show it holds for $t+1$. First, note that $E^{*}_{t}$ cannot decrease so $p_{t}^{\prime}$ cannot increase as $t$ increases. Then, by Equation 14, $p_{t+1}<p_{t+1}^{\prime}$ only if we sample more than $\frac{a_{3}^{\prime}\log^{2}(n)}{2\varepsilon\eta^{2}}$ elements of $E_{\tilde{\alpha}}^{t}$ at probability $p_{t}=p_{t}^{\prime}$ before Algorithm 4.1. However, we have just shown that with high probability,
@@
Now, combining Lemma 4.7 with Lemma 4.9 gives us our final accuracy guarantee in Theorem 4.1. In terms of space complexity, we need $O\left(\frac{\log^{2}(n)\log(\tilde{\alpha})}{\varepsilon\eta^{2}}\right)$ space to store the sampled edges and the two counters per edge.
4.2 Bounded Arboricity Sparsifiers
For our bounded arboricity graphs, we take inspiration from the bounded arboricity sparsifier of Solomon [solomon2021local]. A closely related line of work is that of edge degree constrained subgraphs (EDCS) [assadi2019coresets, bernstein2015fully, bernstein2016faster]. We modify the sparsifiers from [solomon2021local] to show Proposition 4.11, which states that node-neighboring graphs have small edge edit distance post-sparsification.
Our proofs use the sensitivity of the arboricity of the graph which we prove below.
Proposition 4.10 (Node Sensitivity of Arboricity).
The node-sensitivity of the arboricity of a graph $G$ is $1$.

### v3

@@ D.1 Detailed Algorithm Description
The pseudocode for our algorithm is given in Algorithm D.1. Our algorithm works as follows. We set the probability of sampling edges from the stream initially to $1$. For each update $e_{t}$, we sample the update into $S$ with probability $p_{t}$ (Algorithm D.1). If the sampled update is not $\bot$ (Algorithm D.1), then we initialize counters for its endpoints if one or both of these counters do not already exist (Algorithm D.1). A sampled $e_{t}$ that is $\bot$ does not increase the size of $S$ (it is a no-op). Then, we iterate through every other edge in $S$ (Algorithm D.1) and if $e^{\prime}$ shares an endpoint with $e_{t}$ (Algorithm D.1), then we increment the counter $c_{e^{\prime}}^{w}$ associated with $e^{\prime}$ (Algorithm D.1). If the counter exceeds our cutoff $\tilde{\alpha}$ (Algorithm D.1), then we remove $e^{\prime}$ and its corresponding counters from $S$ (Algorithm D.1). After performing our edge sampling, we now check using SVT whether $|S|$ (the number of edges in $S$) exceeds $\frac{a_{3}\log^{2}n}{\varepsilon\eta^{2}}$ for some large enough constant $a_{3}>0$ (Algorithm D.1). If the SVT AboveThreshold query is satisfied, then we halve the probability of sampling (Algorithm D.1) and resample each edge in $S$ with $1/2$ probability (Algorithm D.1). Finally, we output a new estimate if $|S|/p$ is greater than our previous estimate by a significantly large amount. Specifically, we check via SVT queries (Algorithm D.1) whether $|S|$ exceeds $p\cdot(1+\eta)^{j_{t}}$. If so, we increment the counter $j_{t}$ and output our new estimate as $(1+\eta)^{j_{t}}$ (Algorithm D.1); otherwise, we output our old estimate.
@@ 31                         With probability $\nicefrac{{1}}{{2}}$ remove $e$ from $S$ and delete $c_{e}^{u},c_{e}^{v}$
Algorithm D.1 Sampling $E_{t}^{*}$
D.2 Privacy Guarantee
@@
Lemma D.1.
Algorithm D.1 is $\varepsilon$-edge differentially private.
@@
By D.2, since the sensitivity of the sampled set of edges $S$ is upper bounded by 2. The subsample SVT output is $\varepsilon/2$-DP by Theorem 4.10. Now, observe that the probability $p$ of sampling only changes whenever the subsample SVT (Algorithm D.1) accepts. Since privacy is preserved after post-processing, this implies the statement in the claim. ∎
@@
By definition, ${\mathcal{P}}$ and ${\mathcal{P}}^{\prime}$ defines the sampling probability for all elements in ${\mathcal{S}}$ and ${\mathcal{S}}^{\prime}$ to be $p_{t}$. Hence, it holds that $\frac{\mathbf{P}[{\mathcal{I}}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})})|{\mathcal{P}}]}{\mathbf{P}[{\mathcal{I}}^{\prime}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})})|{\mathcal{P}}^{\prime}]}=1$. By D.3, $\frac{\mathbf{P}[{\mathcal{P}}]}{\mathbf{P}[{\mathcal{P}}^{\prime}]}\leq e^{\varepsilon/2}$ and our claim follows. ∎
Let ${\mathcal{Q}}$ be the event that $({\mathcal{I}}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{(T)}),{\mathcal{P}})$ and define ${\mathcal{Q}}^{\prime}$ analogously. {adjustwidth}1em1em
Claim D.5.
Let $\mathbf{a}\in\{\text{``above''},\text{``below''},\text{``abort''}\}^{T}$ be a sequence of the estimate SVT answers in Algorithm D.1. We have
@@
By D.2, since the sensitivity of the sampled set of edges $S$ is upper bounded by 2, the estimate SVT output is $\varepsilon/2$-DP by Theorem 4.10. ∎
@@ D.3 Approximation and Space Guarantees
Let $E^{*}_{t}=\max_{t^{\prime}\leq t}(|E_{\tilde{\alpha}}^{t^{\prime}}|)$, we now show that we get a $\left(1+\eta,\frac{\log^{2}(n)}{\varepsilon\eta}\right)$-approximation of $E_{t}^{*}$ with high probability. First, we note that $E^{*}_{t}$ satisfies $M(G^{t})\leq E^{*}_{t}\leq(2+\tilde{\alpha})M(G_{t})$ since $E^{*}_{t}\geq|E_{\tilde{\alpha}}^{t}|$, $M(G_{t^{\prime}})\leq M(G_{t})$, and Lemma D.6.
@@
Lemma D.8.
Algorithm D.1 returns a $\left(1+\eta,\frac{\log^{2}(n)}{\varepsilon\eta}\right)$-approximation of $E_{t}^{*}$ for every $t\in[T]$ with probability at least $1-1/\operatorname*{poly}(n)$.
@@
First, let $\tau=\frac{d\log^{2}(n)}{\varepsilon\eta^{2}}$ for some large enough constant $d>0$ and define level $i_{t}$ (starting with $i_{t}=2$) to be the level that contains $E_{t}^{*}$ if $2^{i_{t}-1}\cdot\tau<E^{*}_{t}\leq 2^{i_{t}}\cdot\tau$. Level $i_{t}=1$ is defined as $0\leq E^{*}_{t}\leq 2\cdot\tau$. Let $p_{t}$ denote the marginal probability that an edge is sampled from $E_{\alpha}^{t}$. Note that this probability is prior to sub-sampling on Algorithm D.1. In a perfect situation, at time $t$, edge $e$ is sampled in level $i$ with probability $p_{t}^{\prime}:=2^{-i_{t}}$ when $i_{t}\geq 2$ and with probability $p_{t}^{\prime}=1$ when $i_{t}=1$, but this is not the case as our algorithm determines $p_{t}$ adaptively depending on the size of the sampled edges at time $t$.
SVT Budget. We show that with probability at least $1-1/\operatorname*{poly}(n)$, neither of the SVT “above” budgets $Q_{1}$, $Q_{2}$ in Algorithm D.1 are exceeded. First, for the sub-sampling SVT, we take $T=\operatorname*{poly}(n)$ noisy comparisons with ${\textsf{Lap}}(O(\varepsilon/\log(n)))$ noise. Hence with probability $1-1/\operatorname*{poly}(n)$, all Laplace realizations are of order $O(\varepsilon^{-1}\log^{2}(n))$.
$Q_{1}=a_{1}\log(n)$ bounds the number of times the subsampling SVT in Algorithm D.1 is satisfied. Each time this SVT is satisfied, we must have
@@
for sufficiently large constant $a_{3}>0$, leading to the probability of sampling edges being halved. The expected number of edges that are sampled at any point is $p_{t}\cdot\lvert E_{\tilde{\alpha}}^{t}\rvert\leq p_{t}\cdot n^{2}$. Suppose $a_{1}$ is sufficiently large and we halved the sampling probability $\nicefrac{{Q_{1}}}{{2}}\geq\lceil\log(n^{2})\rceil$ times so that $p_{t}\leq\frac{1}{n^{2}}$. We can upper bound the probability that the SVT answers “above” by considering the experiment where we obtain a sample $\tilde{S}_{t}$ by sampling each edge with probability $\tilde{p}=\frac{1}{\lvert E_{\tilde{\alpha}}^{t}\rvert}\geq p_{t}$. By a multiplicative Chernoff bound (Theorem A.1), for $\psi=\frac{a_{3}}{2}\varepsilon^{-1}\log^{2}(n)-1$,
@@
Concentration. In a perfect situation, edge $e$ is sampled at time $t$ with probability $p^{\prime}_{t}=2^{-i_{t}}$ for $i_{t}\geq 2$ and with probability $p^{\prime}_{t}=1$ for $i_{t}=1$. In that case, we can use the multiplicative Chernoff bound to bound the probability that our estimate concentrates as our sampling probability would be sufficiently high enough to do so. However, it is not the case that $p_{t}$ is guaranteed to be $2^{-i_{t}}$ since it is determined adaptively by Algorithm D.1. Hence, we need a slightly more sophisticated analysis than simply using the multiplicative Chernoff bound with $p^{\prime}_{t}$.
We consider two cases. If $p_{t}\geq p^{\prime}_{t}$, then we can lower bound the probability of success by what we would obtain using $p_{t}^{\prime}$ and we can use the multiplicative Chernoff bound in this case. Now, suppose that $p_{t}<p^{\prime}_{t}$; then, we claim that this case never occurs in Algorithm D.1 with probability at least $1-1/\operatorname*{poly}(n)$. We now formally present these arguments.
If $p_{t}\geq p_{t}^{\prime}$, then we use a multiplicative Chernoff bound (Theorem A.1) to bound our probability of success. Let $s_{t}$ be the number of edges we sampled for $E_{\tilde{\alpha}}^{t}$ prior to any sub-sampling. That is, we obtain $s_{t}$ by sampling with probability $p_{t}$ before Algorithm D.1. For $\psi:=\eta\cdot E_{t}^{*}/\lvert E_{\tilde{\alpha}}^{t}\rvert$,
@@
Now, we consider the case when $p_{t}<p_{t}^{\prime}$. We show that with probability at least $1-1/\operatorname*{poly}(n)$, this case does not occur. We prove this via induction on $t$. For $t=1$, $p_{t}=p_{t}^{\prime}$ trivially since both equal $1$. Now, we assume our claim holds for $t$ and show it holds for $t+1$. First, note that $E^{*}_{t}$ cannot decrease so $p_{t}^{\prime}$ cannot increase as $t$ increases. Then, by Equation 109, $p_{t+1}<p_{t+1}^{\prime}$ only if we sample more than $\frac{a_{3}^{\prime}\log^{2}(n)}{2\varepsilon\eta^{2}}$ elements of $E_{\tilde{\alpha}}^{t}$ at probability $p_{t}=p_{t}^{\prime}$ before Algorithm D.1. However, we have just shown that with high probability,
@@
Lemma D.9.
Algorithm D.1 requires $O\left(\frac{\log^{2}(n)\log(\tilde{\alpha})}{\varepsilon\eta^{2}}\right)$ space, with high probability.